OpenStudy (anonymous):

y=10cosx at the point (π/3,5), what is the derivative

OpenStudy (dumbcow):

derivative of cos is -sin --> dy/dx = -10sin x

OpenStudy (anonymous):

how will i find these two also The slope of the tangent line is m1= . The equation of the tangent line is y= .

OpenStudy (dumbcow):

slope = dy/dx evaluated at some x

OpenStudy (anonymous):

im not familiar with that method. would you mind demonstrating it please?

OpenStudy (dumbcow):

what method? its just the rule, derivative of sin is cos derivative of cos is -sin

OpenStudy (anonymous):

cos/-sin is the solution?

OpenStudy (dumbcow):

?? your answer will have 3 parts a) derivative, dy/dx b) slope of tangent line, take dy/dx plug in x=pi/3 c) equation of tangent line, use slope and find y_intercept

OpenStudy (anonymous):

i dont understand how to do any of those. im working ahead on my homework b/c im going to busy with other things. so my teacher hasnt cover those yet and i am clueless. can you help me out please?

OpenStudy (dumbcow):

i already gave you the derivative, now just plug in pi/3 for x to get slope

OpenStudy (anonymous):

-10sin(pi/3) ok

OpenStudy (dumbcow):

yep...now use unit circle or calculator to find out what sin(pi/3) is

OpenStudy (anonymous):

60degrees (1/2,sqrt3/2

OpenStudy (anonymous):

what do i do with those numbers

OpenStudy (dumbcow):

which one refers to sin

OpenStudy (anonymous):

sqrt3/2

OpenStudy (dumbcow):

cos() --> similar to x coord sin() --> similar to y coord good so slope is -5sqrt(3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the line tangent?

OpenStudy (dumbcow):

y = mx+b m = -5sqrt3 (x,y) = (pi/3,5) find b

OpenStudy (anonymous):

wait, i didnt mention that my homework software was satisfied with -10sinsqrt(3) as the slope

OpenStudy (dumbcow):

you mean...-10sin(pi/3)

OpenStudy (anonymous):

haha yeah

OpenStudy (anonymous):

my homework software accepted that

OpenStudy (dumbcow):

ok, same thing...you can interchange sqrt3/2 and sin(pi/3) because they are equal y = mx+b b = y - mx b = 5 -(-10sin(pi/3)*pi/3)

OpenStudy (anonymous):

ok thanks now i have to do slope and tangent line for m2

OpenStudy (anonymous):

wait

OpenStudy (dumbcow):

m2?? haha, same function different point

OpenStudy (anonymous):

my software didnt accept the previous answer

OpenStudy (dumbcow):

what did you put in? did you enter it as y=mx+b form

OpenStudy (anonymous):

i entered it in this form 5 -(-10sin(pi/3)*pi/3)

OpenStudy (dumbcow):

well what does that number represent ?

OpenStudy (dumbcow):

i was trying to help you, i didn't give you the final answer

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

5 -(-10sin(pi/3)*pi/3) i got 5+10sin(pi/3)pi/3

OpenStudy (anonymous):

when i simplify

OpenStudy (dumbcow):

good

OpenStudy (dumbcow):

not done yet...what did you just solve for? what answer are they looking for..."equation of tangent line"

OpenStudy (anonymous):

the line that runs diagonally and only touches a parabola on its sides

OpenStudy (dumbcow):

yes thats definition of tangent line. equation of tangent line is : y = mx+b so answer must be in that form: slope*x + y_intercept

OpenStudy (anonymous):

our slope is 10sin(pi/3) and y-int is 5 right?

OpenStudy (dumbcow):

no

OpenStudy (dumbcow):

y = mx+b b = y - mx m = -10sin(pi/3) (x,y) = (pi/3,5) b = 5 -(-10sin(pi/3)*pi/3)

OpenStudy (anonymous):

i dont see how it isnt. you have m=-10sin(pi/3) and y=5

OpenStudy (dumbcow):

the y_coord of the point is 5. don't confuse that with y_intercept, where line crosses y-axis we represent y_intercept as b

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (anonymous):

im not sure how you solve for b here

OpenStudy (dumbcow):

b = y - mx b = 5 -(-10sin(pi/3)*pi/3)

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

i didnt see the y-mx

OpenStudy (anonymous):

i got 45

OpenStudy (dumbcow):

really? after multiplying by pi....or did you round it to 45

OpenStudy (dumbcow):

i get 14.069

OpenStudy (anonymous):

my calc was in radians mode

OpenStudy (dumbcow):

yeah thats what you want you could probably leave it as: b= 5+10(pi/3)*sin(pi/3),

OpenStudy (anonymous):

ahh i didnt distribute the negative outside the parantheisis

OpenStudy (anonymous):

whats next

OpenStudy (dumbcow):

enter the equation of tangent line y = ?

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=y+%3D+10cos%28x%29+find+tangent+line+at+x%3Dpi%2F3 here is a nice reference to check your answers

OpenStudy (anonymous):

ok thanks m2, do find the derivative of the derivative of m1

OpenStudy (dumbcow):

i dunno what they mean by m2

OpenStudy (dumbcow):

is their another function, or another point ?? does it look like dy^2/dx

OpenStudy (anonymous):

The slope of the normal line is m2 The equation of the normal line is y

OpenStudy (dumbcow):

oh ok, so m1 refers to tangent line m2 refers to normal line which is perpendicular to tangent line at point (pi/3, 5)

OpenStudy (dumbcow):

this will be easier for perpendicular lines: m1*m2 = -1 m2 = -1/m1

OpenStudy (dumbcow):

Then find equation of normal line same way solve for y_intercept b = y - m2*x

OpenStudy (anonymous):

what is that link for?

OpenStudy (anonymous):

thanks for the help

OpenStudy (dumbcow):

your welcome link is to check your answer